OpenStudy (anonymous):

what is the last 4 digit prime ending in "53" starting from 9953, 9853, 9753 etc. ?

OpenStudy (anonymous):

Start with 1053.\[\text{PrimeQ}[1053]\to \text{False} \]\[\text{PrimeQ}[1153]\to \text{True} \]PrimeQ is a Mathematica function that returns True if the argument is prime and False otherwise.

OpenStudy (rational):

@BSwan try this

OpenStudy (rational):

find all \(\large x\) and \(\large y\) such that : \(\large 1000x + 100y+ 53\) is a prime

OpenStudy (anonymous):

when its up to primes , only one thing work -.- guessing hehehe

OpenStudy (anonymous):

yeah lol , why dont you use matlab ?

OpenStudy (rational):

\(\large 1000x + 100y + 53 = p\) \(\large 1000x + 100y = p-53\) this is a linear diophantine equation, solution exist only when \(\large 100 | ( p-53)\)

OpenStudy (anonymous):

hmmm

OpenStudy (rational):

thats an useless thing as it doesn't tell anything new :o

OpenStudy (anonymous):

wait brb

OpenStudy (anonymous):

i dont think diphontane solve anything right ?

OpenStudy (anonymous):

so what about trial ?